# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 5.3.5. Summary of Adversarial Defenses
Li et al. (Li et al., 2022a) 2022 ICSE Adversarial Training DL-CAIS PbNN Code authorship attribution
Ramakrishnan et al. (Henkel et al., 2022) 2022 SANER Adversarial Training Seq2Seq Code2Seq Code summarization
Cristina et al. (Improta et al., 2023) 2023 CoRR Adversarial Training Seq2Seq CodeBERT CodeT5+ Code generation
DaK (Gao et al., 2023) 2023 PACM_CL Adversarial Training CodeBERT Code2Vec GGNN Code summarization Method name prediction Variable misuse prediction
He et al. (He and Vechev, 2023) 2023 CCS Model Modification CodeGen Code generation
Li et al. (Li et al., 2022b) 2022 COLING Additional Models CodeBERT GraphCodeBERT Defect detection Natural language code search Code question answering
Zhang et al. (Zhang et al., 2023e) 2023 arXiv Additional Models GPT-3.5 GPT-4 Claude-Instant-1 Claude-2 Code summarization
Yang et al. (Yang et al., 2024e) 2024 TOSEM Additional Models CodeGPT PLBART CodeT5 Replit CodeGen CodeT5+ Code generation
Adversarial defenses for CodeLMs encompass input modification, model modification, and additional model methods, each with a specific focus on mitigating the effects of adversarial attacks. The input modification approach often involves adversarial training, which has been shown to enhance model robustness. Model modification strategies adapt sub-networks, activation functions, or loss functions to detect and handle adversarial samples effectively. Lastly, the additional model approach leverages external models for classification robustness. Methods like RoPGen, SVEN, and SPACE contribute significantly to adversarial robustness, highlighting the potential of combining adversarial training with program-specific features for enhanced defense against complex adversarial scenarios.
@@ 6. Answer to RQ3: Empirical Studies on Security of CodeLMs
2021 Nguyen (Nguyen et al., 2021) Adversarial Attack 1) How well has the issue of AML in RSSE been addressed by the existing literature? 2) To what extent are state-of-the-art API and code snippet recommender systems susceptible to malicious data?
2023 Du (Du et al., 2023) Adversarial Attack 1) How do existing adversarial attack approaches perform against different PTMCs under various tasks? 2) What is the quality of adversarial examples generated by adversarial attacks? 3) How do the contexts of the perturbed identifiers affect the adversarial attacking performance?
2023 Li (Li et al., 2023a) Adversarial Defense 1) How does learning sample obtaining strategy during training impact model performance? 2) How does loss computation impact model performance? 3) How does input data impact model performance? 4) How does input model impact model performance?
2024 Oh (Oh et al., 2024) Backdoor Attack 1) How do developers’ adoption rates and trust levels differ when using code completion compared to code generation as AI-powered coding assistant tools, and what factors influence these variations? 2) How do poisoning attacks on AI-powered coding assistant tools influence the security of software developers’ code in the real world? 3) Which type of AI-powered coding assistant tools, code completion or code generation , are more susceptible to poisoning attacks?
2024 Hussain (Hussain et al., 2024a) Backdoor Attack 1) Is there any trojan signature in the classifier weights of the trojaned code models which were produced by full-finetuning? 2) Is there any trojan signature in the trojaned code models, which were produced where pretrained weights were frozen?
2024 Hussain (Hussain et al., 2024b) Backdoor Attack 1) Distribution of attention weight and bias in clean and poisoned models. 2) Clustering activation values of clean and poisoned samples. 3) Visualizing context embeddings of clean and poisoned models. 4) Comparison of fine-tuned parameters with pre-trained parameters. 5) Resetting fine-tuned weights to pre-trained weights based on a threshold.
@@ 6.3. Effectiveness and robustness of adversarial training for CodeLMs
Li et al. (Li et al., 2023a) conduct an empirical study on the effectiveness and robustness of adversarial training for protecting CodeLMs. First, they collect and categorize five types of adversarial training methods for CodeLMs: mixing directly, composite loss, adversarial fine-tuning, min-max + composite loss, and min-max. Next, they conduct empirical evaluations of these categorized adversarial training methods on the code summarization task using Code2Seq (Alon et al., 2019a) and the code authorship attribution task using Abuhamad (Abuhamad et al., 2018). Specifically, they investigate the following research questions: 1) How does learning sample obtaining strategy during training impact model performance? Different adversarial training methods may adopt various strategies to acquire the samples to be learned in the current training iteration. For example, perturbing the input samples or selecting samples that maximize the loss for learning. They evaluate these two strategies to study their impact on the performance of CodeLMs. They find that the impact of sample obtaining strategies for CodeLM varies depending on the task, and no single strategy is suitable for all tasks. However, the performance of optimization objective based methods using the min-max strategy is related to the number of times perturbed samples are regenerated. Therefore, they suggest that extending training time can be traded off for higher model robustness. 2) How does loss computation impact model performance? Some adversarial training methods use ordinary loss, where all terms are computed equally for all samples. Other methods use composite loss, where some terms are computed for clean samples and others for adversarial or perturbed samples. They find that composite loss is generally more effective than ordinary loss. However, in terms of model robustness, composite loss performs better when using object-oriented methods, while ordinary loss demonstrates stronger robustness when combined with data augmentation-like methods. 3) How does input data impact model performance? Some adversarial training methods require the input to be adversarial samples, while others require the use of both adversarial and perturbed samples. Even within the same category of samples, the composition may vary. They find that the input data used in adversarial training has a significant impact on the model’s robustness, with adversarial samples being the key factor, rather than perturbed samples. When generating adversarial samples, the type of transformations is highly correlated with the final model’s robustness. They recommend including data with various transformations to cover potential attacks that could reduce robustness. 4) How does input model impact model performance? They evaluate the impact of input models on the final model’s robustness from two aspects: the influence of pre-trained models and the number of fine-tuning epochs. They find that pre-trained models outperform non-pre-trained models in terms of training time and can surpass composite loss methods in robustness, achieving comparable robustness to direct mixing methods. Additionally, for adversarial fine-tuning, only a few fine-tuning epochs are required on pre-trained models, as increasing the number of fine-tuning epochs does not lead to better model performance.
@@ 7. Answer to RQ4: Experimental Setting and Evaluation
7.1. RQ3.1: Experimental Datasets
In this section, we present the experimental datasets used in CodeLM security research.
@@
Dataset Year Programming Language Data Source Download Link Study
BigCloneBench (Svajlenko et al., 2014) 2014 Java GitHub https://github.com/clonebench/BigCloneBench (Yang et al., 2022; Na et al., 2023; Tian et al., 2023; Zhang et al., 2023b; Nguyen et al., 2023; Hussain et al., 2023a; Li et al., 2024b; Awal et al., 2024; Yang et al., 2024a)
OJ dataset (Kim, 2014) 2016 C++ OJ Platform http://programming.grids.cn (Zhang et al., 2020a; Tian et al., 2021; Zhang et al., 2022a; Tian et al., 2023; Zhang et al., 2024d)
CodeSearchNet (Husain et al., 2019) 2019 Go Java JavaScript PHP Python Ruby GitHub https://github.com/github/CodeSearchNet (Li et al., 2022b, 2023b) (Li et al., 2022b, 2023b; Qi et al., 2023; Sun et al., 2023a; Jha and Reddy, 2023; Chen et al., 2022; Yang et al., 2024d) (Li et al., 2022b, 2023b) (Li et al., 2022b, 2023b; Jha and Reddy, 2023) (Li et al., 2022b, 2023b; Yang et al., 2024c; Wan et al., 2022; Choi et al., 2022; Henkel et al., 2022; Qi et al., 2023; Sun et al., 2023a; Jha and Reddy, 2023; Ramakrishnan and Albarghouthi, 2022; Yang et al., 2024d) (Li et al., 2022b, 2023b)
Devign (Zhou et al., 2019) 2019 Java GitHub https://github.com/rjust/defects4j (Zhang et al., 2023b; Hussain et al., 2023a; Li et al., 2024b)
CodeXGLUE (Lu et al., 2021) 2021 Java C/C++ Python PHP JavaScript Ruby Go GitHub https://github.com/microsoft/CodeXGLUE (Li et al., 2023b; Yang et al., 2024d) (Li et al., 2023b; Nguyen et al., 2023) (Li et al., 2023b, b) (Li et al., 2023b) (Li et al., 2023b) (Li et al., 2023b) (Li et al., 2023b)
Shellcode_IA32 (Liguori et al., 2021) 2021 assembly language instruction GitHub https://github.com/dessertlab/Shellcode_IA32 (Improta et al., 2023)
Table 10 provides a detailed list of datasets used in the field of CodeLM security, along with relevant information for each dataset. This information includes the dataset name, references, publication time, programming languages included, data volume, data source, and download address links. This compilation aims to facilitate researchers in downloading and using these datasets for future studies.
@@
7.2. RQ3.2: Experimental LMs for Code
To better understand the behavior and vulnerabilities of CodeLMs, numerous CodeLMs have been developed and applied in the field of program analysis and code-related tasks. In this section, we introduce and summarize the experimental CodeLMs used in adversarial and backdoor research on CodeLM security.
@@
LSTM 1997 - (Quiring et al., 2019; Tian et al., 2021; Zhou et al., 2022; Zhang et al., 2024c, d; Bielik and Vechev, 2020; Zhang et al., 2022a; Yi et al., 2023; Li et al., 2024b)
@@
GCN 2013 - (Zhang et al., 2024d; Bielik and Vechev, 2020)
GRU 2014 - (Zhang et al., 2024d, 2022a)
GGNN 2015 - (Gao et al., 2023; Yu et al., 2023; Zhang et al., 2024d; Yefet et al., 2020)
Transformer 2017 - (Wan et al., 2022; Li et al., 2024b; Qi et al., 2023; Zhang et al., 2023d)
@@
CodeBERT 2020 https://huggingface.co/microsoft/codebert-base (Pour et al., 2021; Choi et al., 2022; Li et al., 2022b; Zhang et al., 2022a; Yang et al., 2022; Jha and Reddy, 2023; Gao et al., 2023; Nguyen et al., 2023; Na et al., 2023) (Improta et al., 2023; Zhang et al., 2023c, 2024c; Tian et al., 2023; Li et al., 2024b; Wan et al., 2022; Sun et al., 2023a; Cotroneo et al., 2024) (Yang et al., 2024c; Qi et al., 2023; Yang et al., 2024a; Zhang et al., 2024d; Zhou et al., 2024; Yang et al., 2024d; Hussain et al., 2023a)
GraphCodeBERT 2020 https://huggingface.co/microsoft/graphcodebert-base (Yang et al., 2022; Li et al., 2022b; Zhang et al., 2024c; Jha and Reddy, 2023; Nguyen et al., 2023; Na et al., 2023; Zhang et al., 2023c, c; Qi et al., 2023; Tian et al., 2023; Yang et al., 2024a; Zhang et al., 2024d; Zhou et al., 2024)
CodeT5 2021 https://huggingface.co/Salesforce/codet5-base (Jha and Reddy, 2023; Na et al., 2023; Zhang et al., 2023c; Yang et al., 2024e; Sun et al., 2023a; Tian et al., 2023; Yang et al., 2024c; Li et al., 2023b; Yang et al., 2024a; Zhou et al., 2024; Yang et al., 2024d; Hussain et al., 2023a)
CodeT5+ 2021 https://huggingface.co/Salesforce/codet5-base (Improta et al., 2023; Cotroneo et al., 2024; Yang et al., 2024e)
PLBART 2021 https://github.com/wasiahmad/PLBART (Yang et al., 2024c; Li et al., 2023b; Yang et al., 2024e; Hussain et al., 2023a)
CodeGen 2022 https://huggingface.co/Salesforce/codegen-350M-multi (He and Vechev, 2023; Yang et al., 2024e, d; Aghakhani et al., 2024)
GPT-3.5 2022 https://openai.com/ (Zhang et al., 2023e)
GPT-4 2023 https://openai.com/ (Zhang et al., 2023e)
@@
This section introduces the evaluation metrics used in attacks and defenses targeting CodeLMs. These metrics are typically employed to measure attack success rates, model robustness, and the effectiveness of defense methods.
@@
Average Normalized Rank ($ANR$).: This metric is used to measure the effectiveness of backdoor attacks against code retrieval models. $ANR$ indicates how much the attack can improve the retrieval ranking of poisoned samples. The smaller $ANR$ value represents the more effective attack method. Specifically, $s^{\prime}$ stands for the length of the injected trigger code snippet. $|S|$ stands for the length of the full ranking list. The calculation formula of $ANR$ is as follows:
@@
3
Number of Queries. This represents the average number of successful queries against the target model in an adversarial attack. For black-box adversarial attack methods, queries are the only way to access the target model. Therefore, the number of queries is one of the important indicators to evaluate the efficiency of black-box adversarial attack methods. Specifically, $q_{i}$ represents the number of queries for the $i_{th}$ successful attack. In detail, $i\in\{j|f(j)=1\}$. The calculation formula of $NumberofQueries$ is as follows:
@@
6
Valid Rate ($V_{r}$). $V_{r}$ is defined as the percentage of adversarial samples that can be compiled (Tramèr et al., 2016). This metric is used to evaluate the quality of the adversarial samples generated quantity as well as the efficiency of the generation process. Specifically, $Count_{valid}$ is the number of adversarial samples that can pass the compile check of the program. $Count_{all}$ is the number of all samples. The calculation formula of $V_{r}$ is as follows:
@@
8
Variable Change Rate ($VCR$). $VCR$ is the proportion of variables that an attacker needs to modify for a successful attack. The lower the VCR, the better the attack method, and the more efficient the adversarial sample can be found only by modifying a small number of variables. In the formula for Variable Change Rate (VCR), $n_{i}$ refers to the number of variables that an attacker has successfully modified in the $i$-th adversarial sample, while $m_{i}$ represents the total number of variables present in the $i$-th sample. The calculation formula of $VCR$ is as follows:
@@
$BP$ stands for brevity penalty. When ${\displaystyle r\leq c}$, the brevity penalty ${\displaystyle BP=1}$, meaning that we do not punish long candidates and only punish short candidates. When ${\displaystyle r>c}$, the brevity penalty ${\displaystyle BP=e^{1-r/c}}$
@@
Attack Success Rate ($ASR$). $ASR$ refers to the proportion of poisoned samples that are successfully predicted by the model to target label, which is frequently used to measure the effect of backdoor attacks. In general, the higher the $ASR$, the better the performance of the attack method. $M$ refers to clean model, $M^{\prime}$ refers to backdoored model. $C$ refers to clean data, $C^{\prime}$ refers to poisoned data. The calculation formula of $ASR$ is as follows:
7

### v2

@@ 5.3.5. Summary of Adversarial Defenses
RoPGen (Li et al., 2022a) 2022 ICSE Input modification DL-CAIS PbNN Code authorship attribution
Henkel et al. (Henkel et al., 2022) 2022 SANER Input modification Seq2Seq Code2Seq Code summarization
CARROTT (Zhang et al., 2022a) 2022 TOSEM Input modification GRU, LSTM ASTNN, LSCNN TBCNN, CodeBERT CDLH Function classification Clone detection Vulnerability detection
Cristina et al. (Improta et al., 2023) 2023 ArXiv Input modification Seq2Seq CodeBERT CodeT5+ Code generation
EverI (Gao et al., 2023) 2023 PLDI Input modification CodeBERT Code2Vec GGNN Code summarization Method name prediction Variable misuse prediction
SVEN (He and Vechev, 2023) 2023 CCS Model modification CodeGen Code generation
SPACE (Li et al., 2022b) 2022 COLING Additional models CodeBERT GraphCodeBERT Defect detection Natural language code search Code question answering
Zhang et al. (Zhang et al., 2023c) 2023 ArXiv Additional models GPT-3.5, GPT-4 Claude-Instant-1 Claude-2 Code summarization
RADAR-Defense (Yang et al., 2024g) 2024 TOSEM Additional models CodeGPT, PLBART CodeT5, CodeT5+ Replit, CodeGen Code generation
Table 9 summarizes the existing adversarial defense methods for CodeLMs. Adversarial defenses for CodeLMs encompass input modification, model modification, and additional model methods, each with a specific focus on mitigating the effects of adversarial attacks. The input modification approach often involves adversarial training, which has been shown to enhance model robustness. Model modification strategies adapt sub-networks, activation functions, or loss functions to detect and handle adversarial samples effectively. Lastly, the additional model approach leverages external models for classification robustness. Methods like RoPGen, SVEN, and SPACE contribute significantly to adversarial robustness, highlighting the potential of combining adversarial training with program-specific features for enhanced defense against complex adversarial scenarios.
@@ 6. Answer to RQ3: Empirical Studies on Security of CodeLMs
2021 Nguyen et al. (Nguyen et al., 2021) Adversarial Attack 1) How well has the issue of AML in RSSE been addressed by the existing literature? 2) To what extent are state-of-the-art API and code snippet recommender systems susceptible to malicious data?
2023 Du et al. (Du et al., 2023) Adversarial Attack 1) How do existing adversarial attack approaches perform against different PTMCs under various tasks? 2) What is the quality of adversarial examples generated by adversarial attacks? 3) How do the contexts of the perturbed identifiers affect the adversarial attacking performance?
2023 Li et al. (Li et al., 2023b) Adversarial Defense 1) How does learning sample obtaining strategy during training impact model performance? 2) How does loss computation impact model performance? 3) How does input data impact model performance? 4) How does input model impact model performance?
2024 Oh et al. (Oh et al., 2024) Backdoor Attack 1) How do developers’ adoption rates and trust levels differ when using code completion compared to code generation as AI-powered coding assistant tools, and what factors influence these variations? 2) How do poisoning attacks on AI-powered coding assistant tools influence the security of software developers’ code in the real world? 3) Which type of AI-powered coding assistant tools, code completion or code generation , are more susceptible to poisoning attacks?
2024 Hussain et al. (Hussain et al., 2024a) Backdoor Attack 1) Is there any trojan signature in the classifier weights of the trojaned code models which were produced by full-finetuning? 2) Is there any trojan signature in the trojaned code models, which were produced where pretrained weights were frozen?
2024 Hussain et al. (Hussain et al., 2024b) Backdoor Attack 1) Distribution of attention weight and bias in clean and poisoned models. 2) Clustering activation values of clean and poisoned samples. 3) Visualizing context embeddings of clean and poisoned models. 4) Comparison of fine-tuned parameters with pre-trained parameters. 5) Resetting fine-tuned weights to pre-trained weights based on a threshold.
@@ 6.3. Effectiveness and robustness of adversarial training for CodeLMs
Li et al. (Li et al., 2023b) conduct an empirical study on the effectiveness and robustness of adversarial training for protecting CodeLMs. First, they collect and categorize five types of adversarial training methods for CodeLMs: mixing directly, composite loss, adversarial fine-tuning, min-max + composite loss, and min-max. Next, they conduct empirical evaluations of these categorized adversarial training methods on the code summarization task using Code2Seq (Alon et al., 2019a) and the code authorship attribution task using Abuhamad (Abuhamad et al., 2018). Specifically, they investigate the following research questions: 1) How does learning sample obtaining strategy during training impact model performance? Different adversarial training methods may adopt various strategies to acquire the samples to be learned in the current training iteration. For example, perturbing the input samples or selecting samples that maximize the loss for learning. They evaluate these two strategies to study their impact on the performance of CodeLMs. They find that the impact of sample obtaining strategies for CodeLM varies depending on the task, and no single strategy is suitable for all tasks. However, the performance of optimization objective based methods using the min-max strategy is related to the number of times perturbed samples are regenerated. Therefore, they suggest that extending training time can be traded off for higher model robustness. 2) How does loss computation impact model performance? Some adversarial training methods use ordinary loss, where all terms are computed equally for all samples. Other methods use composite loss, where some terms are computed for clean samples and others for adversarial or perturbed samples. They find that composite loss is generally more effective than ordinary loss. However, in terms of model robustness, composite loss performs better when using object-oriented methods, while ordinary loss demonstrates stronger robustness when combined with data augmentation-like methods. 3) How does input data impact model performance? Some adversarial training methods require the input to be adversarial samples, while others require the use of both adversarial and perturbed samples. Even within the same category of samples, the composition may vary. They find that the input data used in adversarial training has a significant impact on the model’s robustness, with adversarial samples being the key factor, rather than perturbed samples. When generating adversarial samples, the type of transformations is highly correlated with the final model’s robustness. They recommend including data with various transformations to cover potential attacks that could reduce robustness. 4) How does input model impact model performance? They evaluate the impact of input models on the final model’s robustness from two aspects: the influence of pre-trained models and the number of fine-tuning epochs. They find that pre-trained models outperform non-pre-trained models in terms of training time and can surpass composite loss methods in robustness, achieving comparable robustness to direct mixing methods. Additionally, for adversarial fine-tuning, only a few fine-tuning epochs are required on pre-trained models, as increasing the number of fine-tuning epochs does not lead to better model performance.
@@ 7. Answer to RQ4: Experimental Setting and Evaluation
7.1. RQ4.1: Experimental Datasets
Dataset Year Programming Language Data Source Download Link Study
BigCloneBench (Svajlenko et al., 2014) 2014 Java GitHub https://github.com/clonebench/BigCloneBench (Yang et al., 2022; Na et al., 2023; Tian et al., 2023; Zhang et al., 2023a; Nguyen et al., 2023; Hussain et al., 2023a; Li et al., 2024b; Awal et al., 2024; Yang et al., 2024a)
OJ dataset (Kim, 2014) 2016 C++ OJ Platform http://programming.grids.cn (Zhang et al., 2020a; Tian et al., 2021; Zhang et al., 2022a; Tian et al., 2023; Zhang et al., 2023d)
CodeSearchNet (Husain et al., 2019) 2019 Go Java JavaScript PHP Python Ruby GitHub https://github.com/github/CodeSearchNet (Li et al., 2022b, 2023c) (Li et al., 2022b, 2023c; Qi et al., 2023; Sun et al., 2023; Jha and Reddy, 2023; Chen et al., 2022; Yang et al., 2024f) (Li et al., 2022b, 2023c) (Li et al., 2022b, 2023c; Jha and Reddy, 2023) (Li et al., 2022b, 2023c; Yang et al., 2024c; Wan et al., 2022; Choi et al., 2022; Henkel et al., 2022; Qi et al., 2023; Sun et al., 2023; Jha and Reddy, 2023; Ramakrishnan and Albarghouthi, 2022; Yang et al., 2024f) (Li et al., 2022b, 2023c)
Devign (Zhou et al., 2019) 2019 Java GitHub https://github.com/rjust/defects4j (Zhang et al., 2023a; Hussain et al., 2023a; Li et al., 2024b)
CodeXGLUE (Lu et al., 2021) 2021 Java C/C++ Python PHP JavaScript Ruby Go GitHub https://github.com/microsoft/CodeXGLUE (Li et al., 2023c; Yang et al., 2024f) (Li et al., 2023c; Nguyen et al., 2023) (Li et al., 2023c, c) (Li et al., 2023c) (Li et al., 2023c) (Li et al., 2023c) (Li et al., 2023c)
Shellcode_IA32 (Liguori et al., 2021) 2021 ASM instruction GitHub https://github.com/dessertlab/Shellcode_IA32 (Improta et al., 2023)
In this section, we present the experimental datasets commonly used in CodeLM security research. Table 11 summarizes these datasets along with key information and is intended to support future research by providing convenient access to commonly used benchmark datasets.
@@
7.2. RQ4.2: Experimental CodeLMs
To better understand the behavior and vulnerabilities of CodeLMs, numerous CodeLMs have been developed and applied in the field of program analysis and code-related tasks. In this section, we introduce and summarize the experimental CodeLMs used in adversarial and backdoor research on CodeLM security. Table 12 provides a detailed list of experimental models used in the field of CodeLMs security.
@@
LSTM 1997 - (Quiring et al., 2019; Tian et al., 2021; Zhou et al., 2022; Zhang et al., 2024c, 2023d; Bielik and Vechev, 2020; Zhang et al., 2022a; Yi et al., 2023; Li et al., 2024b)
@@
GCN 2013 - (Zhang et al., 2023d; Bielik and Vechev, 2020)
GRU 2014 - (Zhang et al., 2023d, 2022a)
GGNN 2015 - (Gao et al., 2023; Yu et al., 2023; Zhang et al., 2023d; Yefet et al., 2020)
Transformer 2017 - (Wan et al., 2022; Li et al., 2024b; Qi et al., 2023; Zhang et al., 2023b)
@@
CodeBERT 2020 https://huggingface.co/microsoft/codebert-base (Pour et al., 2021; Choi et al., 2022; Li et al., 2022b; Zhang et al., 2022a; Yang et al., 2022; Jha and Reddy, 2023; Gao et al., 2023; Nguyen et al., 2023; Na et al., 2023) (Improta et al., 2023; Zhang et al., 2023a, 2024c; Tian et al., 2023; Li et al., 2024b; Wan et al., 2022; Sun et al., 2023; Cotroneo et al., 2024) (Yang et al., 2024c; Qi et al., 2023; Yang et al., 2024a; Zhang et al., 2023d; Zhou et al., 2024; Yang et al., 2024f; Hussain et al., 2023a)
GraphCodeBERT 2020 https://huggingface.co/microsoft/graphcodebert-base (Yang et al., 2022; Li et al., 2022b; Zhang et al., 2024c; Jha and Reddy, 2023; Nguyen et al., 2023; Na et al., 2023; Zhang et al., 2023a, a; Qi et al., 2023; Tian et al., 2023; Yang et al., 2024a; Zhang et al., 2023d; Zhou et al., 2024)
CodeT5 2021 https://huggingface.co/Salesforce/codet5-base (Jha and Reddy, 2023; Na et al., 2023; Zhang et al., 2023a; Yang et al., 2024g; Sun et al., 2023; Tian et al., 2023; Yang et al., 2024c; Li et al., 2023c; Yang et al., 2024a; Zhou et al., 2024; Yang et al., 2024f; Hussain et al., 2023a)
CodeT5+ 2021 https://huggingface.co/Salesforce/codet5-base (Improta et al., 2023; Cotroneo et al., 2024; Yang et al., 2024g)
PLBART 2021 https://github.com/wasiahmad/PLBART (Yang et al., 2024c; Li et al., 2023c; Yang et al., 2024g; Hussain et al., 2023a)
CodeGen 2022 https://huggingface.co/Salesforce/codegen-350M-multi (He and Vechev, 2023; Yang et al., 2024g, f; Aghakhani et al., 2024)
GPT-3.5 2022 https://openai.com/ (Zhang et al., 2023c)
GPT-4 2023 https://openai.com/ (Zhang et al., 2023c)
@@
This section introduces the evaluation metrics used in attacks and defenses targeting CodeLMs. These metrics are typically employed to measure attack success rates, model robustness, and the effectiveness of defense methods. It is worth noting that these evaluation metrics are of interest to both attackers and defenders, as the same metrics can be used to assess the effectiveness of attack methods as well as the resilience of defense methods.
@@
Attack Success Rate ($ASR$). $ASR$ refers to the proportion of poisoned samples that are successfully predicted by the model to target label, which is frequently used to measure the effect of backdoor attacks or defenses. In general, the higher the $ASR$, the better the performance of the attack method, and vice versa. $M$ refers to clean model, $M^{\prime}$ refers to backdoored model. $C$ refers to clean data, $C^{\prime}$ refers to poisoned data. The calculation formula of $ASR$ is as follows:
3
Average Normalized Rank ($ANR$). This metric is used to measure the effectiveness of backdoor attacks or defenses on code retrieval models. $ANR$ indicates how much the attack can improve the retrieval ranking of poisoned samples. The smaller $ANR$ value represents the more effective attack method, and vice versa. Specifically, $s^{\prime}$ stands for the length of the injected trigger code snippet. $|S|$ stands for the length of the full ranking list. The calculation formula of $ANR$ is as follows:
@@
Therefore, the number of queries is not only an important indicator for evaluating the efficiency of black-box adversarial attack methods but also a crucial metric for assessing the effectiveness of corresponding adversarial defense methods.
4
Number of Queries. This represents the average number of successful queries against the target model in an adversarial attack. For black-box adversarial attack methods, queries are the only way to access the target model. Therefore, the number of queries is not only an important indicators for evaluating the efficiency of black-box adversarial attack methods , but also an important metric for assessing the effectiveness of corresponding adversarial defense methods. Specifically, $q_{i}$ represents the number of queries for the $i_{th}$ successful attack. In detail, $i\in\{j|f(j)=1\}$. The calculation formula of Number of Queries is as follows:
@@
7
Valid Rate ($V_{r}$). $V_{r}$ is defined as the percentage of adversarial samples that can be compiled (Tramèr et al., 2016). This metric is used to evaluate the quality of the adversarial samples generated during adversarial attacks or adversarial defenses, as well as the efficiency of the generation process. Specifically, $Count_{valid}$ is the number of adversarial samples that can pass the compile check of the program. $Count_{all}$ is the number of all samples. The calculation formula of $V_{r}$ is as follows:
@@
9
Variable Change Rate ($VCR$). $VCR$ is the proportion of variables that an attacker needs to modify for a successful attack. The lower the VCR, the better the attack method, and the more efficient the adversarial sample can be found only by modifying a small number of variables. Conversely, the higher the $VCR$, the more effective the applied defense method. In the formula for $VCR$, $n_{i}$ refers to the number of variables that an attacker has successfully modified in the $i$-th adversarial sample, while $m_{i}$ represents the total number of variables present in the $i$-th sample. The calculation formula of $VCR$ is as follows:
@@
$BP$ stands for brevity penalty. When ${\displaystyle r\leq c}$, the brevity penalty ${\displaystyle BP=1}$, meaning that we do not punish long candidates and only punish short candidates. When ${\displaystyle r>c}$, the brevity penalty ${\displaystyle BP=e^{1-r/c}}$.